OpenStudy (anonymous):

What is one example each of how the Fibonacci sequence and the Golden Ratio are found in nature? Explain your examples.

OpenStudy (anonymous):

i for some reason think the fibonacci serious was found by observing rabbits mate... If i remember right

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

huh????

OpenStudy (anonymous):

hey jimmyrep, you look like my old math teacher haha

OpenStudy (anonymous):

i dont think i remember it well enough to explain it... but you have 1 pair of rabbits who mate, first cycle thye have another pair of rabbits but there is still only 1 pair... next cycle they have another pair, but the first offspring can mate now. next cycle three pairs can make, next cycle five p[airs can make 1,1,2,3,5,8,13,........

OpenStudy (anonymous):

still only one pair that can mate*

OpenStudy (anonymous):

http://www.youtube.com/watch?v=14-NdQwKz9w watch that, it has to do with your question and quite entertaining

OpenStudy (anonymous):

the golden ratio is involved in nuclear magnetc resonance spin equations for some cobalt compounds

OpenStudy (anonymous):

can you explain the whole golden ratio example but simpler please?

OpenStudy (anonymous):

i read something very amusing about the golden ratio and human anatomy (don't know if its true) - the height of a human divided by the height of his navel = golden ratio!!

OpenStudy (anonymous):

it = 1.618 to 3 DP's

OpenStudy (anonymous):

im so confused. THe guy gave a good example on the fibanic or whatever, but i need an example like that for the golden.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/File:SimilarGoldenRectangles.svg (a+b)/a golden ratio

OpenStudy (anonymous):

OMG im so confused..you gave a perfect explanation for fib. what about golden?

OpenStudy (anonymous):

How about the ratio of the length of the forearm to the length of the hand? Basically, the golden ratio says (a+b)/a = a/b. So in the arm/hand example, the ratio of the length of the arm (from elbow to fingertips) over the forearm length (from elbow to wrist) is the same as the length of the hand (from wrist to finger tips) over the forearm length (from elbow to wrist.) Elbow to wrist = a wrist to fingertips = b Elbow to fingertips = a + b (a+b)/a = a/b